Mongas205:

What value of x is in the solution set of –2(3x + 2) > –8x + 6? –6 –5 5 6

Mongas205:

just a question no diagrams

supie:

Do you know the first step?

Mongas205:

no

supie:

Ok so the first step is to distribute `-2` to everything inside of the parenthesis which is `3x` and `2` So we're doing \(-2(3x)=?\\-2(2)=?\)

Mongas205:

ok

supie:

Do you know? \(-2(3x)=?\\-2(2)=?\)

Mongas205:

no

supie:

for \(-2(3x)=?\) simply do -2 x 3 then add on the x to the end-

Mongas205:

6x

supie:

-6x but close enough.

Mongas205:

ok

supie:

So then for the second one just do the same thing just without x

supie:

-2 x 2

Mongas205:

-4

supie:

Yes so now we have \(−6x−4>−8x+6\) we get `-8x` and change the minus `(-)` to a plus `(+)` so now that it's a plus we have `+8x` now we add 8x to both sides of the equation \(−6x−4+8x>−8x+6+8x\)

supie:

So I believe we should get \(2x−4>6\) then we do that same method again get `-4` change the minus to a plus now we have +4 so we add 4 to both sides of the equation we should get \(2x>10\) because \(2x−4+4>6+4\) 6+4=10

supie:

Do you know what we do next?

Mongas205:

idk

supie:

oof you divide both sides by 2 you think you can do that?

supie:

Ok so we'll have \(\LARGE\frac{2x}{2}=\frac{10}{2}\) \(\cancel{\frac{2x}{2}}=\frac{10}{2}\) (2x/2 cancels) all you have left is 10/2 so simplify that then you should have your answer \(\frac{10}{2}=?\) x=?

Mongas205:

5

supie:

Yes so \(x>5\)

Mongas205:

yayyyy

Mongas205:

thank you

supie:

np :)

Mongas205:

do you have a website to help

supie:

Help with?

Mongas205:

math

supie:

I don't really know of any.

Mongas205:

ok thank you anyways

Mongas205:

o

Mongas205:

no im good thank you

Mongas205:

im just writing the steps

TogaHimiko:

i saw my bad srry fot botherin

Mongas205:

its ok love

TogaHimiko:

mk

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Ok so we'll have \(\LARGE\frac{2x}{2}=\frac{10}{2}\) \(\cancel{\frac{2x}{2}}=\frac{10}{2}\) (2x/2 cancels) all you have left is 10/2 so simplify that then you should have your answer \(\frac{10}{2}=?\) x=? \(\color{#0cbb34}{\text{End of Quote}}\) good job - perfect explained - like a math teacher - CONGRATS !!! @supie

supie:

Ty so much